# Comparative Evaluation of Acoustic Feature Extraction Tools for Clinical Speech Analysis

### v1

@@ Abstract
This study compares three acoustic feature extraction toolkits—OpenSMILE, Praat, and Librosa—applied to clinical speech data from individuals with schizophrenia spectrum disorders (SSD) and healthy controls (HC). By standardizing extraction parameters across the toolkits, we analyzed speech samples from 77 SSD and 87 HC participants and found significant toolkit-dependent variations. While F0 percentiles showed high cross-toolkit correlation (r=0.962–0.999), measures like F0 standard deviation and formant values often had poor, even negative, agreement. Additionally, correlation patterns differed between SSD and HC groups. Classification analysis identified F0 mean, HNR, and MFCC1 (AUC >0.70) as promising discriminators. These findings underscore reproducibility concerns and advocate for standardized protocols, multi-toolkit cross-validation, and transparent reporting.
@@ 3.1 Data collection
Participants diagnosed with schizophrenia spectrum disorders (SSD; N = 77, females = 24.7%, mean age = 35.92) were enrolled from both inpatient and outpatient departments at a hospital in the US. Participants underwent screening using the psychosis and mood modules of the Structured Clinical Interview for DSM-IV [21] and were confirmed to meet DSM-5 diagnostic criteria for schizophrenia spectrum disorders. Healthy volunteers (N = 87, females = 51.7%, mean age = 36.12) were also enrolled either through their previous involvement in other research studies or by responding to online advertisements. All participants gave informed consent, with minors providing assent. All study procedures were approved by the Institutional Review Board. All participants performed several speech-based tasks, including three picture description tasks, as part of a larger study. All speech samples were digitally recorded, and later manually time-stamped and annotated by trained human annotators.
@@ 3.4 Acoustic Features and their Clinical Relevance
Mel-frequency cepstral coefficients (MFCCs) provide a compact representation of the speech spectrum that roughly corresponds to the auditory system’s response. The first MFCC (MFCC1) primarily reflects overall spectral energy distribution, while higher coefficients (MFCC2-4) capture increasingly fine spectral details related to vocal tract configuration and articulation. In schizophrenia research, altered MFCC patterns have been associated with both production differences (potentially related to articulatory precision or vocal tract tension) and perceptual characteristics that listeners identify as “disorganized” speech [39, 40].
The clinical utility of these acoustic features extends beyond simple group discrimination to monitoring treatment effects, predicting functional outcomes, and potentially serving as objective biomarkers. However, this potential can only be realized if the features can be extracted reliably and consistently across different software implementations—the central focus of our present investigation.
@@ 3.5 Statistical analysis
The analysis involved several statistical approaches to evaluate consistency and reliability across the three toolkits. Pearson correlation coefficients were calculated between each toolkit pair for every acoustic feature to assess agreement. Statistical significance of correlations was determined using standard t-tests with p-value thresholds of 0.05, 0.01, and 0.001. To evaluate whether correlations differed significantly between toolkit pairs, Fisher’s r-to-z transformation was applied, comparing correlation coefficients with rigorous statistical testing. Additionally, the classification potential of features for distinguishing SSD from HC groups was assessed using ROC curve analysis and calculating Area Under Curve (AUC) values, with features showing AUC values above 0.7 considered to have good discrimination potential.
@@ 4 Results
Figure 1: Correlation heatmap comparing acoustic feature extraction across three tools (OpenSMILE [OS], Praat [PR], and Librosa [LR]) for SSD and HC groups. Color intensity indicates correlation strength from -1 (dark blue) to 1 (dark red). Statistical significance of correlations is marked: ns (not significant), * (p<0.05), ** (p<0.01), and *** (p<0.001). Empty cells indicate toolkit pairs not available in specific features.

### v2

@@ Abstract
This study compares three acoustic feature extraction toolkits—OpenSMILE, Praat, and Librosa—applied to clinical speech data from individuals with schizophrenia spectrum disorders (SSD) and healthy controls (HC). By standardizing extraction parameters across the toolkits, we analyzed speech samples from 33 SSD and 38 HC participants and found significant toolkit-dependent variations. While F0 percentiles showed high cross-toolkit correlation (r=0.962–0.999), measures like F0 standard deviation and formant values often had poor, even negative, agreement. Additionally, correlation patterns differed between SSD and HC groups. Classification analysis identified F0 mean, HNR, and MFCC1 (AUC >0.70) as promising discriminators. These findings underscore reproducibility concerns and advocate for standardized protocols, multi-toolkit cross-validation, and transparent reporting.
@@ 3.1 Data collection
Participants diagnosed with schizophrenia spectrum disorders (SSD; N = 33, females = 24.7%, mean age = 35.92) were enrolled from both inpatient and outpatient departments at a hospital in the US. Participants underwent screening using the psychosis and mood modules of the Structured Clinical Interview for DSM-IV [21] and were confirmed to meet DSM-5 diagnostic criteria for schizophrenia spectrum disorders. Healthy volunteers (N = 38, females = 51.7%, mean age = 36.12) were also enrolled either through their previous involvement in other research studies or by responding to online advertisements. All participants gave informed consent, with minors providing assent. All study procedures were approved by the Institutional Review Board. All participants performed several speech-based tasks, including three picture description tasks, as part of a larger study. All speech samples were digitally recorded, and later manually time-stamped and annotated by trained human annotators.
@@ 3.4 Acoustic Features and their Clinical Relevance
Mel-frequency cepstral coefficients (MFCCs) provide a compact representation of the speech spectrum that roughly corresponds to the auditory system's response. The first MFCC (MFCC1) primarily reflects overall spectral energy distribution, while higher coefficients (MFCC2-4) capture increasingly fine spectral details related to vocal tract configuration and articulation. In schizophrenia research, altered MFCC patterns have been associated with both production differences (potentially related to articulatory precision or vocal tract tension) and perceptual characteristics that listeners identify as ``disorganized'' speech [39, 40].
The clinical utility of these acoustic features extends beyond simple group discrimination to monitoring treatment effects, predicting functional outcomes, and potentially serving as objective biomarkers. However, this potential can only be realized if the features can be extracted reliably and consistently across different software implementations – the central focus of our present investigation.
@@ 3.5 Statistical analysis
The analysis involved several statistical approaches to evaluate consistency and reliability across the three toolkits. Pearson correlation coefficients were calculated between each toolkit pair for every acoustic feature to assess agreement. Statistical significance of correlations was determined using standard t-tests with p-value thresholds of 0.05, 0.01, and 0.001. To evaluate whether correlations differed significantly between toolkit pairs, Fisher's r-to-z transformation was applied, comparing correlation coefficients with rigorous statistical testing. Additionally, the classification potential of features for distinguishing SSD from HC groups was assessed using ROC curve analysis and calculating Area Under Curve (AUC) values, with features showing AUC values above 0.7 considered to have good discrimination potential.
@@ 4 Results
Figure 1: Correlation heatmap comparing acoustic feature extraction across three tools (OpenSMILE [OS], Praat [PR], and Librosa [LR]) for SSD and HC groups. Color intensity indicates correlation strength from -1 (dark blue) to 1 (dark red). Statistical significance of correlations is marked: * (p<0.05), ** (p<0.01), and *** (p<0.001). Empty cells indicate toolkit pairs not available in specific features.